If z = x + iy and if the point P in the Argand plane represents z, find the locus of z satisfying
the equation |z-2-3i|=5

Grade:upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
given equation is|z-2-3i|=5 where z=x+iy
So |x-2+i(y-3)|=5
(x-2)2+(y-3)2=25
x2+y2-4x-6y-12=0 is the locus
Thanks and Regards
Shaik Aasif
askIITians faculty